Solve: (logbase16 x)(logbase16 x)=1/4
Answer by stanbon(75887)

Solve: (logbase16 x)(logbase16 x)=1/4
-------------------------------------------
(logbase16 x)^2 - 1/4 =0
Factor to get:
(log16 x -1/2)(log16 x + 1/2 ) = 0
log16 x = 1/2 or log16 x = -1/2
x = 16^(1/2) or x = 16^(-1/2)
x = 4 or x = 1/4
=====================
cheers,
Stan H.